100ml of PH3 when decomposed produced phosporous and hydrogen. The change in volume is

Grade:12th pass

1 Answers

Arun
6 years ago
The equation for the decomposition is:
                      
  4PH3 (g) = P4 (s)+ 6H2 (g)  (Elemental phosphorus exists as tetraatomic molecules).

When the temperature and pressure remain the same, 4 ml of PH3 will give 6 ml of H2.   (On applying Gay-Lussac's law ).

Hence, 100 ml of PH3 will yield 150 ml of H2.

Increase in volume = (150-100) ml = 50 ml. (The volume of solid phosphorus is negligible).

There is a 50% increase in volume.